# Negative feedback and oscillations in a model for mRNA translation

### v1

@@ Abstract
The ribosome flow model (RFM) is a phenomenological model for the unidirectional flow of particles along a 1D chain of $n$ sites. The RFM has been extensively used to study the dynamics of ribosome flow along a single mRNA molecule during translation. In this case, the particles model ribosomes and each site corresponds to a consecutive group of codons. Networks of interconnected RFMs have been used to model and analyze large-scale translation in the cell and, in particular, the effects of competition for shared resources. Here, we analyze the RFM with a negative feedback connection from the protein production rate to the initiation rate. This models, for example, the production of proteins that inhibit the translation of their own mRNA. Using tools from the theory of 2-cooperative dynamical systems, we provide a simple condition guaranteeing that the closed-loop system admits at least one non-trivial periodic solution. When this condition holds, we also explicitly characterize a large set of initial conditions such that any solution emanating from this set converges to a non-trivial periodic solution. Such a solution corresponds to a periodic pattern of ribosome densities along the mRNA, and to a periodic pattern of protein production.
@@
mRNA translation, periodic solutions, regulation of gene expression.
@@ I Introduction
Computational modeling of ribosome flow is attracting considerable interest [44, 7, 51]. This is motivated also by new experimental methods, like ribosome profiling, that provide considerable information on which mRNAs are being actively translated in a cell at a given time. The totally asymmetric simple exclusion process (TASEP) [4, 23, 37] describes the flow of particles in $n$ sites along a 1D lattice. Each site may be empty or contain a particle, and particles hop randomly in a unidirectional manner, that is, from site $i$ to site $i+1$. Importantly, a particle can only hop into an empty site (simple exclusion), yielding an intricate coupling between the particles. In particular, if a particle is “stuck” at a site for an extended time, then other particles will accumulate in the sites behind it, thus generating a “traffic jam”. TASEP has found many useful applications, including modeling the flow of ribosomes along mRNA [49] and other intracellular transport processes [5], pedestrian and vehicular traffic [14], the movement of ants, and numerous other natural and artificial phenomena [37]. However, rigorous analysis of TASEP in the case where the transition rates between sites are different is non-trivial. Existing results, for example, on contraction in TASEP [10, 11] do not provide the level of detail that deterministic models can provide.
@@
In the case where the entry, exit, and elongation rates of the RFM are time-varying and jointly periodic, with a common period $T$, the dynamics admit a unique $T$-periodic solution that is globally exponentially stable [25]. Interpreting the rates as a periodic excitation (e.g., due to the periodic cell-cycle process) implies that the state variables in the RFM, and thus also the protein production rate, entrain to the periodic excitation (see also [36] for some related considerations).
Gene expression is tightly regulated, and negative feedback loops are common motifs that can drive the protein production rate to a desired level, as well as reduce expression noise and intercellular variability in protein levels [38]. For example, bacteria use RNA binding proteins (RBPs) as regulators of many cellular processes, including translation, and most RBPs affect translation at the level of translation initiation. Many translation-related RBPs inhibit their own mRNA [16, 2]. Note also that in the context of mRNA transcription, it is known that a negative feedback mechanism induces oscillations in the notch effector Hes1 [15].
Here, we analyze for the first time the RFM with a negative feedback interconnection from the protein production rate to the initiation rate. We show that:
@@
    The closed-loop system admits a unique equilibrium point $e$. This corresponds to a dynamic steady-state where the ribosome flow into and out of each site are equal, and thus to a constant protein production rate;
@@
    Under a simple and easy to verify sufficient condition on the local (in)stability of the system at $e$, the closed-loop system admits a periodic solution, and we provide an explicit description of a set $S$ such that any solution emanating from $S$ converges to a periodic solution.
These results suggest that negative feedback in translation can provide a simple mechanism for producing self-sustained (i.e., unforced) oscillations in protein production. It is known that translational control of cyclin B1 mRNA might play an essential role in the periodic cell-cycle process [9]. This is attributed to oscillations in the eIF4E binding factor, whose expression is cell cycle regulated. However, our results suggest that the simpler mechanism of negative feedback in translation can also lead to periodic oscillations in protein production. This agrees with the findings in [42] that used ribosome profiling to conclude that gene-specific translational repression controls the mitotic proteome, complementing post-translational mechanisms for inactivating protein function.
The fact that negative feedback may lead to oscillations is not surprising, but rigorously proving this in $n$-dimensional non-linear systems like the RFM is non-trivial. Note also that property 2) above implies that the closed-loop system is not contractive nor cooperative. Our analysis is based on showing that the closed-loop system is a strongly 2-cooperative system [46, 21].
The remainder of this paper is organized as follows. The next section describes the RFM with negative feedback. Section III describes the main analysis results. The last section concludes. The proofs of the main results are placed in the Appendix.
@@ II The model
where $x_{i}(t)\in[0,1]$ is the ribosome density levels at site $i$ at time $t$, and $\lambda_{i}>0,i=0,...,n$ are the translation rates from site $i$ to site $i+1$ at time $t$. To explain this non-linear model, consider the general equation
@@
This can be explained as follows. The term $\lambda_{i-1}x_{i-1}(1-x_{i})$ is the flow rate from site $i-1$ to site $i$. This depends on: (1) the parameter $\lambda_{i-1}$ that models “how easy it is” for a ribosome to move from site $i-1$ to site $i$ (e.g., due to the abundance of the cognate tRNA molecules); (2) the density $x_{i-1}$ at site $i-1$; and (3) the “free space” $1-x_{i}$ at site $i$. This is a “soft” version of the simple exclusion principle: as the density in site $i$ increases, the flow into this site decreases. Similarly, the second term on the right-hand side of (2) is the flow rate from site $i$ to site $i+1$.
@@
with $k>0$ and $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a continuously differentiable function such that $g^{\prime}(z)=\frac{dg(z)}{dz}<0$ for any $z>0$. Thus, $g$ is a monotonically decreasing function. We refer to $k$ as the feedback gain. This models negative feedback from the protein production rate $y(t)$ to the RFM initiation rate (that is controlled by the input $u(t)$).
@@ III Main results
We can now state our main results. The proofs of these results are given in the Appendix. Our first result guarantees in particular that the unit cube $[0,1]^{n}$ is a forward invariant set for the closed-loop system. In other words, if we initialize the state-variables $x_{i}(0)$ at time zero with admissible density values, that is, all values are in the range $[0,1]$ then they maintain admissible values for all $t>0$.
@@
If $z\in\mathbb{R}^{n}$ has no zero entries then the number of sign variation in $z$, denoted $\sigma(z)$, is the number of indexes $i$ such that $z_{i}z_{i+1}<0$. For example, for $z=\begin{bmatrix}1&-2&4\end{bmatrix}^{\top}$, we have $z_{1}z_{2}<0$ and $z_{2}z_{3}<0$, so $\sigma(z)=2$. More generally, for any vector $z\in\mathbb{R}^{n}$, with $z\not=0$, the weak number of sign variations in $z$ is defined as $s^{-}(z):=\sigma(\tilde{z})$, where $\tilde{z}$ is the vector obtained from $z$ after deleting all its zero entries. For example, for $z=\begin{bmatrix}0&1&-2&0&4\end{bmatrix}^{\top}$, we have $\tilde{z}=\begin{bmatrix}1&-2&&4\end{bmatrix}^{\top}$, so $s^{-}(z)=\sigma(\tilde{z})=2$. For the zero vector, define $s^{-}(0):=0$. Note that for any $z\in\mathbb{R}^{n}$ and any non-zero scalar $c$, we have $s^{-}(cz)=s^{-}(z)$.
@@
Let $e\in(0,1)^{n}$ denote the unique equilibrium of the closed-loop system described by (II) and (4). If $J(e)=\frac{\partial}{\partial x}f(e)$ has at least two unstable eigenvalues then the following property holds. For any initial condition $a\in[0,1]^{n}$ such that $a\not=e$ and $s^{-}(a-e)\leq 1$ the corresponding solution $x(t,a)$ converges to a non-trivial periodic orbit.
Note that since $e\in(0,1)^{n}$, we have $s^{-}(0-e)=0$ implying that if $J(e)$ has at least two unstable eigenvalues then the solution $x(t,0)$ always converges to a non-trivial periodic orbit. The zero initial condition corresponds in our model to no ribosomes along the mRNA molecule, which is the case for a newly synthesized mRNA.
@@
Consider the closed-loop system with $n=3$, $\lambda_{0}=12$, $\lambda_{1}=6$, $\lambda_{2}=5$, $\lambda_{3}=8$, and $u(t)=kg(y(t))$, where $g(s)=\frac{1}{1+s^{20}}$. When $k=0.1$, a calculation shows that $e=\begin{bmatrix}0.1855&0.2004&0.1113\end{bmatrix}^{\top}$ (all numerical values are to 4-digit accuracy), and the eigenvalues of $J(e)$ are
$-14.9362,\;-2.7564+6.0039i,\;-2.7564-6.0039i.$
@@
Figure 1: Several trajectories of the closed-loop system in Example 1 with $k=0.1$. The initial conditions are marked with $\circ$ and the equilibrium point $e$ is marked by *.
Figure 2: A trajectory of the closed-loop system in Example 1 with $k=1$. The equilibrium point $e$ is marked by *.
@@ IV Discussion
More work is needed in order to analyze the properties of the oscillatory solutions in our model e.g. their amplitude and frequency, and on how these depend on the various parameters of the system.

### v2

@@ Abstract
The ribosome flow model (RFM) is a phenomenological model for the unidirectional flow of particles along a 1D chain of $n$ sites. The RFM has been extensively used to study the dynamics of ribosome flow along a single mRNA molecule during translation. In this case, the particles model ribosomes and each site corresponds to a consecutive group of codons. Networks of interconnected RFMs have been used to model and analyze large-scale translation in the cell and, in particular, the effects of competition for shared resources. Here, we analyze the RFM with a negative feedback connection from the protein production rate to the initiation rate. This models, for example, the production of proteins that inhibit the translation of their own mRNA. The RFM with negative feedback is a 2-cooperative dynamic system, i.e. its flow maps the set of vectors with up to one sign variation to itself. Using tools from the theory of 2-cooperative dynamical systems, we provide a simple condition guaranteeing that the closed-loop system admits at least one non-trivial periodic solution. When this condition holds, we also explicitly characterize a large set of initial conditions such that any solution emanating from this set converges to a non-trivial periodic solution. Such a solution corresponds to a periodic pattern of ribosome densities along the mRNA, and to a periodic pattern of protein production.
@@
mRNA translation, periodic gene expression, 2-cooperative systems.
@@ I Introduction
I-A Modeling ribosome flow
Computational modeling of ribosome flow is attracting considerable interest [54, 11, 61]. This is motivated also by new experimental methods, like ribosome profiling [23], that provide considerable information on which mRNAs are being actively translated in a cell at a given time.
The totally asymmetric simple exclusion process (TASEP) [8, 31, 47] describes the flow of particles along a 1D lattice with $n$ sites. Each site may be empty or contain a particle, and particles hop randomly in a unidirectional manner, that is, from site $i$ to site $i+1$. Importantly, a particle can only hop into an empty site (simple exclusion), yielding an intricate indirect coupling between the particles. In particular, if a particle is “stuck” at a site for an extended period of time, then other particles will accumulate in the sites behind it, thus generating a “traffic jam”. TASEP has found many useful applications, including modeling the flow of ribosomes along mRNA [59] and other intracellular transport processes [9], pedestrian and vehicular traffic [19], the movement of ants, and numerous other natural and artificial phenomena [47]. However, rigorous analysis of TASEP in the case where the transition rates between sites are different is non-trivial. Existing results, for example, on contraction in TASEP [15, 16] do not provide the level of detail that deterministic models can provide.
@@ I-A Modeling ribosome flow
In the case where the entry, exit, and elongation rates in the RFM are time-varying and jointly periodic, with a common period $T$, the dynamics admit a unique $T$-periodic solution that is globally exponentially stable [34]. Interpreting the rates as a periodic excitation (e.g., due to the periodic cell-cycle process) implies that the state variables in the RFM, and thus also the protein production rate, entrain to the periodic excitation (see also [46] for some related considerations).
Gene expression is tightly regulated, and negative feedback loops are common motifs that can drive the protein production rate to a desired level, as well as reduce expression noise and intercellular variability in protein levels [48].
Gene expression is highly regulated to ensure that the right proteins are made in the right places within the cell. Here, we are interested in analyzing the effects of negative feedback in the level of translation. For example, bacteria use RNA binding proteins (RBPs) as regulators of many cellular processes, including translation, and most RBPs affect translation at the level of translation initiation. Many translation-related RBPs inhibit their own mRNA [21, 3].
To study this process, we analyze for the first time the RFM with a negative feedback interconnection from the protein production rate to the initiation rate. We show that:
@@
    The closed-loop system admits a unique equilibrium point $e$. This corresponds to a steady-state where the ribosome flow into and out of each site are equal, and thus to a constant protein production rate;
@@
    Under a simple and easy to verify sufficient condition on the local (in)stability of the system at $e$, the closed-loop system admits a periodic solution, and we provide an explicit description of a set $S$ such that any solution emanating from $S$ converges to a (non-trivial) periodic solution.
These results suggest that negative feedback in the level of translation only can provide a simple cellular mechanism for producing self-sustained (i.e., unforced) oscillations in protein production. This agrees with the findings in [52] that used ribosome profiling to conclude that gene-specific translational repression controls the mitotic proteome, complementing post-translational mechanisms for inactivating protein function.
The fact that negative feedback may lead to oscillations is not surprising, but rigorously proving this in $n$-dimensional non-linear systems like the RFM is non-trivial. Note also that property 2) above implies that the closed-loop system is not contractive nor cooperative. Our analysis is based on showing that the closed-loop system is a strongly 2-cooperative system [56, 29]. Roughly speaking, strongly 2-cooperative dynamic systems map the set of vectors with up to one sign variation to itself. Such systems have a Poincaré-Bendixson property, which implies that bounded solutions which do not converge to an equilibrium converge to a periodic orbit. Ref. [29] derived specific conditions on the Jacobian at the equilibrium such that periodic orbits exist, and in this case also described a set of initial conditions that lead to a periodic solution.
I-B Negative feedback and cellular oscillations
The emergence of sustained oscillations (via convergence to periodic orbits) in high-dimensional nonlinear dynamical systems is a highly non-trivial question with important applications in systems biology, including the understanding of bio-molecular oscillators ruling cell life-cycle and metabolism, as well as circadian rhythms in hormone secretion, body temperature and metabolic functions [14].
Many cellular oscillatory processes, including certain circadian oscillators, are based on the transcription-translation feedback loop (TTFL), that includes both positive and negative feedback loops [22]. The recurring motifs of negative feedback loops, positive feedback loops, and reciprocal regulation play a crucial role in timing and regulating cell-cycle oscillations, and it is known that sufficiently strong negative feedback induces oscillations [13]. For a very readable account of the role of feedback loops in biological oscillators, see [1, Chapter 6].
The p53–MDM2 loop is a well-known negative feedback circuit that plays a crucial role in regulating cell fate decisions. p53 activates the transcription of many genes, including mdm2. MDM2 inhibits the transcriptional activity of p53 and increases degradation of p53. This mechanism yields oscillations or pulses in p53 levels in response to DNA damage [4, 7, 6].
A negative feedback mechanism in transcription induces oscillations in the notch effector Hes1 [20]. A “minimal” mathematical model for this phenomena, that includes two ordinary differential equations and a time delay in the negative feedback loop, was derived in [27].
An important goal in synthetic biology is building genetic oscillators that can be used for correctly timing and synchronizing various processes. Elowitz and Leibler [32] designed an artificial oscillator in E. coli consisting of a loop with three transcriptional repressors. Negative feedback loops are known to be a key design principle for synthetic biological oscillators [40], but constructing reliable biomolecular oscillators with tunable amplitude and phase is still an open challenge.
The model that we explore here suggests that oscillations can also be generated by direct negative feedback from the produced protein to the mRNA initiation rate.
The remainder of this paper is organized as follows. The next section describes the RFM with negative feedback. Section III describes the main analysis results. The last section concludes. The proofs of the main results are placed in the Appendix 1.
@@ II The model
where $x_{i}(t)\in[0,1]$ is the ribosome density levels at site $i$ at time $t$, and $\lambda_{i}>0,i=0,...,n$ are the translation rates from site $i$ to site $i+1$ at time $t$. Note that the $x_{i}$s are dimensionless and the $\lambda_{i}$s have units of $1/\text{time}$.
To explain this non-linear model, consider the general equation
@@
This can be explained as follows. The term $\lambda_{i-1}x_{i-1}(1-x_{i})$ is the flow rate from site i1i-1 to site $i$. This depends on: (1) the parameter $\lambda_{i-1}$ that models “how easy it is” for a ribosome to move from site i1i-1 to site $i$ (e.g., due to the abundance of the cognate tRNA molecules); (2) the density $x_{i-1}$ at site i1i-1; and (3) the “free space” $1-x_{i}$ at site $i$. This is a “soft” version of the simple exclusion principle: as the density in site $i$ increases, the flow into this site decreases. Similarly, the second term on the right-hand side of (2) is the flow rate from site $i$ to site $i+1$.
@@
with $k>0$ and $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a continuously differentiable function such that $g^{\prime}(z)=\frac{dg(z)}{dz}<0$ for any $z>0$. Thus, $g$ is a monotonically decreasing function. We refer to $k$ as the feedback gain. This models negative feedback directly from the protein production rate $y(t)$ to the RFM initiation rate that is controlled by the input $u(t)$ (see Fig. 1).
Note that (4) captures many negative feedback functions used in models from systems biology and in particular gene regulation. For example, g(y)=ayg(y)=-ay, with $a>0$ (negative linear feedback), $g(y)=e^{-ay}$, with $a>0$ (exponential decay), $g(y)=\frac{1}{a+y}$, with $a>0$ (Michaelis-Menten inhibition), $g(y)=\frac{1}{1+(y/a)^{\ell}}$, with $a,\ell>0$ (Hill type inhibition), $g(y)=1-\tanh(ay)$, with $a>0$, and more.
[…]
Figure 1: RFMIO with feedback $u=kg(y)$.
@@ III Main results
We can now state our main results. The proofs of these results are given in Appendix 1. Our first result guarantees in particular that the unit cube $0,1]^{n}$ is a forward invariant set for the closed-loop system. In other words, if we initialize the state-variables $x_{i}(0)$ at time zero with admissible density values, that is, all values are in the range $[0,1]$ then they maintain admissible values for all $t>0$.
@@
[…]
@@
Let $e\in(0,1)^{n}$ denote the unique equilibrium of the closed-loop system described by (II) and (4). If $J(e)=\frac{\partial}{\partial x}f(e)$ has at least two unstable eigenvalues then the following property holds. For any initial condition $x(0)\in[0,1]^{n}$ such that $x(0)\not=e$ and $s^{-}(x(0)-e)\leq 1$ the corresponding solution $x(t,x(0))$ converges to a non-trivial periodic orbit.
Note that since $e\in(0,1)^{n}$, we have $s^{-}(0-e)=0$ implying that if $J(e)$ has at least two unstable eigenvalues then the solution $x(t,0)$ always converges to a non-trivial periodic orbit. The zero initial condition corresponds in our model to zero ribosomes along the mRNA molecule at the initial time, which is the case for a newly synthesized mRNA.
The next example demonstrates Theorem 3 using a numerical simulation. The Matlab code for generating this simulation is provided in Appendix 2.
@@
Consider the closed-loop system with $n=3$, $\lambda_{0}=12$, $\lambda_{1}=6$, $\lambda_{2}=5$, $\lambda_{3}=8$, and $u(t)=kg(y(t))$, where $g(s)=\frac{1}{1+s^{20}}$. When $k=0.1$, a calculation shows that $e=\begin{bmatrix}0.1855&0.2004&0.1113\end{bmatrix}^{\top}$ (all numerical values in this paper are to 4-digit accuracy), and the eigenvalues of $J(e)$ are
14.9362,2.7564+6.0039i,2.75646.0039i.-14.9362,\;-2.7564+6.0039i,\;-2.7564-6.0039i.
@@
Figure 2: Convergence of several trajectories of the closed-loop system in Example 1 to the equilibrium point with $k=0.1$. The initial conditions are marked with $\circ$ and the equilibrium point $e$ is marked by *.
Figure 3: Convergence of a trajectory of the closed-loop system in Example 1 to a periodic orbit with $k=1$.
The equilibrium point $e$ is marked by *.
@@
Our theoretical results can be easily applied to RFMs with transition rates that are based on biological data. The analysis flow in such a case is demonstrated in the following example. The Matlab code for the following calculations is provided in Appendix 3.
Example 2.
Consider the RFM derived in [57] for the S. cerevisiae gene YBL025W that encodes the protein RRN10, which is related to regulation of RNA polymerase I. This gene has 145 codons (excluding the stop codon), and the corresponding RFM is based on dividing the mRNA into 6 consecutive pieces that include 24, 25, 25, 25, 25, and 21 non-overlapping codons, respectively. The first piece corresponds to the initiation stage. The corresponding RFM thus includes $n=5$ sites and six rates $\lambda_{0},\dots,\lambda_{5}$. These rates were estimated using ribo-seq data (see [57]) yielding
$\begin{bmatrix}\lambda_{0}&\dots&\lambda_{5}\end{bmatrix}=\begin{bmatrix}0.1678&0.2572&0.2758&0.2514&0.2612&0.3002\end{bmatrix}$, (6)
with units of $\text{sec}^{-1}$. Assume that we close the loop around the corresponding RFMIO with the negative feedback function
$g(y)=1-\tanh(ay),$
with $a>0$. This function takes values in $[0,1]$, and for large values of $a$ it decays quickly with $y$, resembling an on-off switch. Note that $g^{\prime}(y)=-\frac{a}{\cosh^{2}(ay)}$.
The equilibrium point $e$ satisfies the equations
$\lambda_{0}k(1-\tanh(a\lambda_{5}e_{5}))(1-e_{1})$ $=\lambda_{1}e_{1}(1-e_{2})$
$=\lambda_{2}e_{2}(1-e_{3})$
$=\lambda_{3}e_{3}(1-e_{4})$
$=\lambda_{4}e_{4}(1-e_{5})$
$=\lambda_{5}e_{5},$
with the $\lambda_{i}$s in (6). This set of equations can be solved numerically for any $k,a$, and Proposition 2 guarantees the existence of a unique solution $e\in(0,1)^{5}$. For example, for $k=10$ and $a=100$ the unique solution is
$e=\begin{bmatrix}0.1038&0.0978&0.1065&0.1003&0.0803\end{bmatrix}^{\top}$. (7)
The Jacobian of the closed-loop system at $e$ is
$J(e)=\left[\begin{smallmatrix}-\lambda_{0}kg(\lambda_{5}e_{5})-\lambda_{1}(1-e_{2})&\lambda_{1}e_{1}&0&0&\lambda_{0}\lambda_{5}(1-e_{1})kg^{\prime}(\lambda_{5}e_{5})\\ \lambda_{1}(1-e_{2})&-\lambda_{1}e_{1}-\lambda_{2}(1-e_{3})&\lambda_{2}e_{2}&0&0\\ 0&\lambda_{2}(1-e_{3})&-\lambda_{2}e_{2}-\lambda_{3}(1-e_{4})&\lambda_{3}e_{3}&0\\ 0&0&\lambda_{3}(1-e_{4})&-\lambda_{3}e_{3}-\lambda_{4}(1-e_{5})&\lambda_{4}e_{4}\\ 0&0&0&\lambda_{4}(1-e_{5})&-\lambda_{4}e_{4}-\lambda_{5}\end{smallmatrix}\right].$
Substituting the numerical values above and calculating the eigenvalues of $J(e)$ gives
0.0112+0.1894i, 0.01120.1894i,0.3855+0.3072i,0.38550.3072i,0.6301.0.0112+0.1894i,\;0.0112-0.1894i,\;-0.3855+0.3072i,\;-0.3855-0.3072i,\;-0.6301.
Thus, two eigenvalues have a positive real part and our theoretical results imply that for any initial condition $x(0)\in[0,1]^{5}$ such that $x(0)\not=e$ and $s^{-}(x(0)-e)\leq 1$ the corresponding solution $x(t,x(0))$ converges to a non-trivial periodic orbit. Fig. 4 depicts the solution $x(t,x(0))$ of the closed-loop system emanating from the initial condition $x(0)=0$. Note that $s^{-}(x(0)-e)=0$. It may be seen that this solution does not converge to $e$, but rather converges to a periodic orbit. Note that after convergence to the periodic pattern, it is straightforward to obtain from the numerical simulation features such as the amplitude and frequency of each density and of the protein production rate.
Figure 4: The state-variables $x_{1}(t)$ (dotted), $x_{2}(t)$ (dashed), and $x_{3}(t)$ (solid) as a function of $t$ in Example 2. Here $t$ is in sec. To avoid cluttering, $x_{4}(t)$ and $x_{5}(t)$ are not plotted.
@@ IV Discussion
More work is needed in order to analyze the properties of the oscillatory solutions in our model e.g. their amplitude and frequency, and on how these features depend on the various parameters of the system.
Our analysis is based on the fact that the RFM with negative feedback is a $2$-cooperative system. It is natural to ask if this structure is preserved when we add more bio-physical features into the RFM. We describe two possible generalizations that preserve the $2$-cooperative structure. First, adding a decay term to each density equation in the RFM, that is, changing (2) to
$\dot{x}_{i}(t)=-c_{i}x_{i}(t)+\lambda_{i-1}x_{i-1}(t)(1-x_{i}(t))-\lambda_{i}x_{i}(t)(1-x_{i+1}(t)),$
with $c_{i}\in\mathbb{R}$. Indeed, such a modification changes the diagonal terms of the Jacobian of the system, but the diagonal terms play no role in determining $2$-cooperativity (see Eq. (10) below). Of course, the $c_{i}$s may change the location of the equilibrium $e$, as well as the spectral properties of $J(e)$.
Another and perhaps more interesting possibility that preserves $2$-cooperativity is the feedback connection of two RFMIOs, as depicted in Fig. 5, with $g_{1}$ a monotonically decreasing function, and $g_{2}$ a monotonically increasing function. Such a closed-loop network may be used as a model for the interaction between two mRNAs, via the proteins that they produce, and our results allow to determine when oscillations appear in such a network.
[…]
Figure 5: Feedback interconnection of two RFMIOs. One RFMIO has dimension $n$ and rates $\lambda_{i}$, and the second has dimension $m$ and rates $\eta_{i}$. They are interconnected via two feedback functions $g_{1}$ and $g_{2}$.